### v4

# The reflection complexity of sequences over finite alphabets

Abstract
In combinatorics on words, the well-studied factor complexity function $\rho_{\mathbf{x}}$ of a sequence $\mathbf{x}$ over a finite alphabet counts, for every nonnegative integer $n$, the number of distinct length-$n$ factors of $\mathbf{x}$. In this paper, we introduce the reflection complexity function $r_{\mathbf{x}}$ to enumerate the factors occurring in a sequence $\mathbf{x}$, up to reversing the order of symbols in a word. We prove a number of results about the growth properties of $r_{\mathbf{x}}$ and its relationship with other complexity functions. We also prove a Morse–Hedlund-type result characterizing eventually periodic sequences in terms of their reflection complexity, and we deduce a characterization of Sturmian sequences. We investigate the reflection complexity of quasi-Sturmian, episturmian, $(s+1)$-dimensional billiard, complementation-symmetric Rote, and rich sequences. Furthermore, we prove that if $\mathbf{x}$ is $k$-automatic, then $r_{\mathbf{x}}$ is computably $k$-regular, and we use the software Walnut to evaluate the reflection complexity of some automatic sequences, such as the Thue–Morse sequence. We note that there are still many unanswered questions about this reflection measure.
Jean-Paul Allouche    John M. Campbell    Shuo Li
Jeffrey Shallit    Manon Stipulanti
Keywords: factor complexity, reflection complexity, reversal, automatic sequence, Sturmian sequence, quasi-Sturmian sequence, episturmian sequence, billiard sequence, Rote sequence, rich sequence, Morse-Hedlund theorem, Walnut.
MSC: Primary 68R15; Secondary 11B85.
1 Introduction
Given an infinite sequence $\mathbf{x}$ over a finite alphabet, it is natural to study the combinatorial properties of the factors of $\mathbf{x}$. (The term factor refers to a contiguous block occurring in $\mathbf{x}$.) For example, writing $\mathbb{N}_{0}=\{0,1,\ldots\}$ and =$\mathbb{N}=\{1,2,\ldots\}$, many authors have studied the factor complexity function $\rho_{\mathbf{x}}\colon\mathbb{N}_{0}\to\mathbb{N}$, which maps $n\geq 0$ to the number of distinct factors of $\mathbf{x}$ of length $n$. Note that $\rho_{\mathbf{x}}(0)=1$, since every—finite or infinite—sequence has a unique factor of length 00; namely, the empty word.
Variations on this definition can be considered as a measure of how “complicated” a sequence is. For example, the abelian complexity function of $\mathbf{x}$ counts the number of factors of $\mathbf{x}$ of a given length, where two factors $u$ and $v$ are considered the same if they have the same length and one is a permutation of the other. Similarly, the cyclic complexity function $c_{\mathbf{x}}$, introduced in 2017 [32], is equal to the number of length-$n$ factors of $\mathbf{x}$, up to equivalence under rotations (cyclic permutations). By analogy, the abelian and cyclic complexity functions lead us to introduce, in this paper, a reflection complexity function on sequences involving reversals.
In addition to the factor, abelian, and cyclic complexity functions indicated above, there have been many different complexity functions on sequences that have been previously introduced. In this regard, we highlight the following in alphabetical order: additive complexity [9], arithmetical complexity [12], gapped binomial complexity [84], $k$-abelian complexity [60], $k$-binomial complexity [83], Kolmogorov complexity [61], Lempel–Ziv complexity [63], Lie complexity [17], linear complexity [75], maximal pattern complexity [59], maximum order complexity [48], opacity complexity [8], open and closed complexity [77], palindrome complexity [4], periodicity complexity [70], privileged complexity [79], relational factor complexity [35], (initial) (non-)repetitive complexity [28, 71], span and leftmost complexity [33], string attractor profile complexity (implicitly defined in [88]; also see [33]), and window complexity [34]. Also see the references in the surveys in [1, 50, 51, 52]. Our reflection complexity function $r_{\mathbf{x}}$, defined below, does not seem to have been previously studied, but may be thought of as natural in terms of its relationships with automatic sequences such as the Thue–Morse sequence. To begin with, we require the equivalence relation $\sim_{r}$ defined below.
Definition 1.
Let $m,n$ be nonnegative integers. Given a finite word $u=u(1)u(2)\cdots u(m)$, its reversal is the word $u^{R}=u(m)u(m-1)\cdots u(1)$, i.e., $u^{R}(i)=u(m+1-i)$ for all $i\in\{1,\ldots,m\}$. A palindrome is a word that is equal to its reversal. Two finite words $u$ and $v$ are reflectively equivalent if $v=u$ or $v=u^{R}$. We denote this equivalence relation by $u\sim_{r}v$.
Example 2.
Over the alphabet $\{\text{{\tt a}},\text{{\tt b}},\ldots,\text{{\tt z}}\}$, the English word reward is reflectively equivalent to drawer, while deed, kayak, and level are palindromes.
Definition 3.
Let $\mathbf{x}$ be a sequence. The reflection complexity function $r_{\mathbf{x}}\colon\mathbb{N}_{0}\to\mathbb{N}$ of $\mathbf{x}$ maps every $n\geq 0$ to the number of distinct length-$n$ factors of $\mathbf{x}$, up to equivalence by $\sim_{r}$.
Example 4.
Let
[…]
denote the Thue–Morse sequence, where the $n$th term in (1) for $n\geq 1$ is defined as the number of $1$’s, modulo $2$, in the base-2 expansion of $n-1$. The initial terms of the integer sequence $(r_{\mathbf{t}}(n))_{n\geq 0}$ are such that
$(r_{\mathbf{t}}(n))_{n\geq 0}=1,2,3,4,6,6,10,10,13,12,16,16,20,20,22,\ldots.$ (2)
We see that $r_{\mathbf{t}}(2)=3$, for example, since there are $3$ length-2 factors of $\mathbf{t}$, up to reflection complexity, i.e., the factors $00$ and $11$ and one member of the equivalence class $\{01,10\}$, with respect to $\sim_{r}$.
The integer sequence in (2) was not, prior to this paper, included in the On-Line Encyclopedia of Integer Sequences [92], which suggests that our notion of “reflection complexity” is new. (Now it is present as sequence A373700.) Also see the work of Krawchuk and Rampersad in [62], which introduced the notion of cyclic/reversal complexity for sequences. The evaluation of reflection complexity functions is closely related to the work of Rampersad and Shallit [80], who investigated sequences $\mathbf{x}$ such that all sufficiently long factors $w$ have the property that $w^{R}$ is not a factor of $\mathbf{x}$. Also, the evaluation of reflection complexities for sequences is related to the enumeration of palindromes contained in sequences; see, e.g., Fici and Zamboni [53].
This paper is organized as follows. In Section 2, we introduce the notation and definitions needed for the paper. In Section 3, we give general results about reflection complexity. In particular, we investigate its growth properties and relationships with other complexity functions. In Section 4, we give a graph-theoretic interpretation of reflection-equivalent classes and prove an inequality for reflection complexity. In Sections 5, 6, and 7, respectively, we investigate reflection complexity for eventually periodic sequences, Sturmian sequences and generalizations, and reversal-closed and rich sequences. Next, in Section 8, we focus on classical automatic sequences and, with the use of the free software Walnut, we prove that the reflection complexity function for automatic sequences is a regular sequence. We also study reflection complexity for famous automatic sequences such as the Thue–Morse sequence. Finally, some further research directions and open questions are considered in Section 9.
2 Preliminaries
Generalities. For a general reference on words, we cite [65]. An alphabet is a finite set of elements called letters. A word over an alphabet $A$ is a finite sequence of letters from $A$. The length of a word, denoted between vertical bars, is the number of its letters (counting multiplicities). The empty word is the only 00-length word, denoted by $\varepsilon$. For all $n\geq 0$, we let $A^{n}$ denote the set of all length-$n$ words over $A$. We let $A^{*}$ denote the set of words over $A$, including the empty word and equipped with the concatenation operation. In order to distinguish finite words and infinite sequences, we write the latter in bold. Except for complexity functions, we start indexing words and sequences at $1$, unless otherwise specified. A factor of a word or a sequence is one of its (finite and contiguous) subblocks. A prefix (resp., suffix) is a starting (resp., ending) factor. Given a word $w$, its $n$th term is written $w(n)$ for $1\leq n\leq|w|$. The factor starting at position $n$ and ending at position $m$ with $1\leq m\leq n\leq|w|$ is written $w[m..n]$. We let $\operatorname{Fac}_{w}$ denote the set of all factors of $w$ and, for each natural number $n$, we let $\operatorname{Fac}_{w}(n)$ denote the set of all length-$n$ factors of $w$. A factor $u$ of a word $w$ over $A$ is right (resp., left) special if $ua$ and $ub$ (resp., $au$ and $bu$) are factors of $w$ for some distinct letters $a,b\in A$. A sequence $\mathbf{x}$ is reversal-closed if, for every factor $w$ of $\mathbf{x}$, the word $w^{R}$ is also a factor of $\mathbf{x}$. A sequence $\mathbf{x}$ is eventually periodic if there exist finite words $u,v$, with $v$ nonempty, such that $\mathbf{x}=uv^{\omega}$ where $v^{\omega}=vvv\cdots$ denotes the infinite concatenation of $v$. A sequence that is not eventually periodic is said to be aperiodic. A sequence is said to be recurrent if every factor occurs infinitely many times; it is uniformly recurrent if each factor occurs with bounded gaps, i.e., for all factors $w$, there is some length $m=m(w)$ such that $w$ occurs in every length-$m$ block.
Morphisms. Let $A$ and $B$ be finite alphabets. A morphism $f\colon A^{*}\to B^{*}$ is a map satisfying $f(uv)=f(u)f(v)$ for all $u,v\in A^{*}$. In particular, $f(\varepsilon)=\varepsilon$, and $f$ is entirely determined by the images of the letters in $A$. For an integer $k\geq 2$, a morphism is $k$-uniform if it maps each letter to a length-$k$ word. A $1$-uniform morphism is called a coding. A sequence $\mathbf{x}$ is pure morphic if it is a fixed point of a morphism, i.e., there exist a morphism $f\colon A^{*}\to A^{*}$ and a letter $a\in A$ such that $\mathbf{x}=f^{\omega}(a)$, where $f^{\omega}(a)=\lim_{n\to\infty}f^{n}(a)$. A sequence $\mathbf{y}$ is morphic if there exist a pure morphic sequence $\mathbf{x}$ and a coding $g\colon A^{*}\to B^{*}$ such that $\mathbf{y}=g(\mathbf{x})$. We let $E\colon\{0,1\}^{*}\to\{0,1\}^{*}$ be the exchange morphism defined by $E(0)=1$ and $E(1)=0$. We naturally extend $E$ to sequences.
Numeration systems. Let $U=(U(n))_{n\geq 0}$ be an increasing sequence of integers with $U(0)=1$. Any integer $n$ can be decomposed in a greedy way as $n=\sum_{i=0}^{t}c(i)\,U(i)$ with non-negative integer coefficients $c(i)$. The word $c(t)\cdots c(0)\in\mathbb{N}^{*}$ is said to be the (greedy) $U$-representation of $n$. By convention, the greedy representation of 00 is the empty word $\varepsilon$, and the greedy representation of $n>0$ starts with a non-zero digit. For $c_{t}\cdots c_{0}\in\mathbb{N}^{*}$, we let $\operatorname{val}_{U}(c(t)\cdots c(0))$ denote the integer $\sum_{i=0}^{t}c(i)\,U(i)$. A sequence $U$ satisfying all the above conditions defines a positional numeration system.
Automatic and regular sequences. For the case of integer base numeration systems, a classical reference on automatic sequences is [7], while [82, 89] treat the case of more exotic numeration systems.
Let $U=(U(n))_{n\geq 0}$ be an positional numeration system. A sequence $\mathbf{x}$ is $U$-automatic if there exists a deterministic finite automaton with output (DFAO) $\mathcal{A}$ such that, for all $n\geq 0$, the $n$th term $\mathbf{x}(n)$ of $\mathbf{x}$ is given by the output $\mathcal{A}(\operatorname{rep}_{U}(n))$ of $\mathcal{A}$. In particular, if $U$ is the sequence of consecutive powers of an integer $k\geq 2$, then $\mathbf{x}$ is said to be $k$-automatic.
It is known that a sequence is $k$-automatic if and only if it is the image, under a coding, of a fixed point of a $k$-uniform morphism [7].
A generalization of automatic sequences to infinite alphabets is the following [7, 82, 89]. Let $U=(U(n))_{n\geq 0}$ be a positional numeration system. A sequence $\mathbf{x}$ is $U$-regular if there exist a column vector $\lambda$, a row vector $\gamma$ and matrix-valued morphism $\mu$ such that $\mathbf{x}(n)=\lambda\mu(\operatorname{rep}_{U}(n))\gamma$. Such a system of matrices forms a linear representation of $\mathbf{x}$. In particular, if $U$ is the sequence of consecutive powers of an integer $k\geq 2$, then $\mathbf{x}$ is said to be $k$-regular.
Another definition of $k$-regular sequences is the following one [7]. Consider a sequence $\mathbf{x}$ and an integer $k\geq 2$. The $k$-kernel of $\mathbf{x}$ is the set of subsequences of the form $(\mathbf{x}(k^{e}n+r))_{n\geq 0}$ where $r\in\{1,2,\ldots,k^{e}\}$. A sequence is $k$-regular if the $\mathbb{Z}$-module generated by its $k$-kernel is finitely generated. A sequence is then $k$-automatic if and only if its $k$-kernel is finite [7].
Sturmian sequences. A sequence $\mathbf{x}$ is Sturmian if its factor complexity function satisfies $\rho_{\mathbf{x}}(n)=n+1$ (see, e.g., [7, 66]). Sturmian sequences have minimal factor complexity among all non-eventually periodic sequences, as proved by Morse and Hedlund [72].
Theorem 5 ([72]).
Let $\mathbf{x}$ be a sequence and let $\ell$ be the number of distinct letters occurring in $\mathbf{x}$. The following properties are equivalent.
  • (a)
    The sequence $\mathbf{x}$ is eventually periodic.
  • (b)
    We have $\rho_{\mathbf{x}}(n)=\rho_{\mathbf{x}}(n+1)$ for some $n\geq 0$.
  • (c)
    We have $\rho_{\mathbf{x}}(n)<n+\ell-1$ for some $n\geq 1$.
  • (d)
    The factor complexity $\rho_{\mathbf{x}}$ is bounded.
Remark 6.
This theorem implies in particular that $\rho_{\mathbf{x}}(n)$ is either bounded or it satisfies $\rho_{\mathbf{x}}(n)\geq n+\ell-1$ for all $n$. Thus the minimal factor complexity among all non-eventually periodic sequences is $\rho_{\mathbf{x}}(n)=n+1$ for all $n$, i.e., the complexity of Sturmian sequences. Actually there is another “growth gap” for $\rho_{\mathbf{x}}$. Recall that a sequence $\mathbf{x}$ is called quasi-Sturmian if there exists a constant $C$ such that, for $n$ large enough, one has $\rho_{\mathbf{x}}(n)=n+C$ (see [31]; also see [37]). It is known that if $\mathbf{x}$ is neither eventually periodic nor quasi-Sturmian, then $\rho_{\mathbf{x}}(n)-n$ tends to infinity: this result is due to Coven [40, Lemma 1.3]; also see Cassaigne’s proof cited in [3, Proof of Theorem 3, p. 23]. Thus
  • (a)
    either $\rho_{\mathbf{x}}(n)$ is bounded, which happens if and only if $\mathbf{x}$ is eventually periodic;
  • (b)
    or else $\rho_{\mathbf{x}}(n)=n+C$ for some constant $C$ and all $n$ large enough, which means that $\mathbf{x}$ is quasi-Sturmian;
  • (c)
    or else $\rho_{\mathbf{x}}(n)-n$ tends to infinity.
One more point (once explained to the first author by Jean Berstel) is that if $\rho_{\mathbf{x}}(n)=n+1$ for $n$ large enough, then $\rho_{\mathbf{x}}(n)=n+1$ for all $n$. Namely, let $n_{0}$ be the least integer $n$ for which $\rho_{\mathbf{x}}(n)=n+1$, and suppose that $n_{0}>1$. Hence $\rho_{\mathbf{x}}(n_{0}-1)\neq n_{0}$. The sequence $\mathbf{x}$ cannot be eventually periodic, since its factor complexity is not bounded. Thus, one has $\rho_{\mathbf{x}}(n)\geq n+1$ for all $n$. Hence, in particular, $\rho_{\mathbf{x}}(n_{0}-1)\geq n_{0}$. Thus $\rho_{\mathbf{x}}(n_{0}-1)>n_{0}$. Since $\rho_{\mathbf{x}}$ is non-decreasing, we have that $n_{0}<\rho_{\mathbf{x}}(n_{0}-1)\leq\rho_{\mathbf{x}}(n_{0})=n_{0}+1$. This gives $\rho_{\mathbf{x}}(n_{0}-1)=n_{0}+1=\rho_{\mathbf{x}}(n_{0})$, which is impossible since $\mathbf{x}$ is not eventually periodic. In other words, in the second item above, if $C=1$, then $\mathbf{x}$ is Sturmian.
3 General results
Given a sequence $\mathbf{x}$, we can decompose its factor complexity function $\rho_{\mathbf{x}}$ and its reflection complexity function $r_{\mathbf{x}}$ by using the following functions: for all $n\geq 0$, we let
  • (a)
    $\operatorname{Unr}_{\mathbf{x}}(n)$ denote the number of “unreflected” length-$n$ factors $w$ of $\mathbf{x}$ such that $w^{R}$ is not a factor of $\mathbf{x}$;
  • (b)
    $\operatorname{Ref}_{\mathbf{x}}(n)$ denote the number of “reflected” length-$n$ factors $w$ of $\mathbf{x}$ such that $w^{R}$ is also a factor of $\mathbf{x}$; and
  • (c)
    $\operatorname{Pal}_{\mathbf{x}}(n)$ denote the number of length-$n$ palindrome factors $w$ of $\mathbf{x}$ (i.e., the palindrome complexity function of $\mathbf{x}$ [4]).
In particular, we have
$\begin{array}[]{l}\rho_{\mathbf{x}}=\operatorname{Unr}_{\mathbf{x}}+\operatorname{Ref}_{\mathbf{x}},\\ r_{\mathbf{x}}=\operatorname{Unr}_{\mathbf{x}}+\dfrac{1}{2}(\operatorname{Ref}_{\mathbf{x}}-\operatorname{Pal}_{\mathbf{x}})+\operatorname{Pal}_{\mathbf{x}}.\end{array}$ (5)
Example 7.
Let $\mathbf{f}=01001010\cdots$ denote the Fibonacci sequence, which is the fixed point of $0\mapsto 01,1\mapsto 0$. Its length-$5$ factors are $u_{1}=01001$, $u_{2}=10010$, $u_{3}=00101$, $u_{4}=01010$, $u_{5}=10100$, and $u_{6}=00100$. Observe that $u_{1},u_{2},u_{3},u_{5}$ are reflected (second type), and $u_{4}$ and $u_{6}$ are palindromes (second and third types). Furthermore $\mathbf{f}$ has no unreflected factor, because Sturmian words are reversal-closed; see [47, Theorem 4, p. 77]. We therefore obtain $r_{\mathbf{f}}(6)=0+\frac{1}{2}(6-2)+2=4$.
The relationships among the complexity functions $\operatorname{Unr}_{\mathbf{x}}$, $\operatorname{Ref}_{\mathbf{x}}$, $\rho_{\mathbf{x}}$, $r_{\mathbf{x}}$, and $\operatorname{Pal}_{\mathbf{x}}$ motivates the study of the combinations of these functions indicated in Equalities (5). This is illustrated below.
Lemma 8.
For a sequence $\mathbf{x}$ and for all $n\geq 1$, we have
$\rho_{\mathbf{x}}(n)-r_{\mathbf{x}}(n)=\frac{1}{2}(\operatorname{Ref}_{\mathbf{x}}(n)-\operatorname{Pal}_{\mathbf{x}}(n))$
and
$2r_{\mathbf{x}}(n)-\rho_{\mathbf{x}}(n)=\operatorname{Unr}_{\mathbf{x}}(n)+\operatorname{Pal}_{\mathbf{x}}(n).$
Proof.
Immediate from Equalities (5). ∎
This lemma implies the following bounds on the ratio $r/\rho$.
Theorem 9.
For a sequence $\mathbf{x}$ and for all $n\geq 0$, we have
$\frac{1}{2}\rho_{\mathbf{x}}(n)\leq\frac{1}{2}(\rho_{\mathbf{x}}(n)+\operatorname{Pal}_{\mathbf{x}}(n))\leq r_{\mathbf{x}}(n)\leq\rho_{\mathbf{x}}(n).$
Furthermore, the equality cases are as follows:
  • (a)
    We have $r_{\mathbf{x}}(n)=\rho_{\mathbf{x}}(n)$ if and only if every reflected length-$n$ factor of $\mathbf{x}$ is a palindrome.
  • (b)
    We have $r_{\mathbf{x}}(n)=\frac{1}{2}(\rho_{\mathbf{x}}(n)+\operatorname{Pal}_{\mathbf{x}}(n))$ if and only if $\mathbf{x}$ has no unreflected length-$n$ factors. In particular, if the sequence $\mathbf{x}$ is reversal-closed, we have $r_{\mathbf{x}}=\frac{1}{2}(\rho_{\mathbf{x}}+\operatorname{Pal}_{\mathbf{x}})$.
  • (c)
    We have $r_{\mathbf{x}}(n)=\frac{1}{2}\rho_{\mathbf{x}}(n)$ if and only if $\mathbf{x}$ has no palindrome of length $n$ and each of its length-$n$ factors is reflected.
Proof.
The inequalities and the equality cases are immediate consequences of Lemma 8. ∎
Remark 10.
It is known that if a sequence $\mathbf{x}$ is reversal-closed, then $\mathbf{x}$ is recurrent: it suffices to adapt the proof of [43, Proposition 1, p. 176], as indicated in [22]. Also note that if a sequence $\mathbf{x}$ is uniformly recurrent and contains infinitely many distinct palindromes, then $\mathbf{x}$ is reversal-closed [15, Theorem 3.2].
One can say more for uniformly recurrent sequences. The following dichotomy holds.
Theorem 11.
Let $\mathbf{x}$ be a uniformly recurrent sequence. Then either it is reversal-closed, or else it has no long reflected factors (which implies that $\mathbf{x}$ has no long palindromes). In other words,
  • (a)
    either $\rho_{\mathbf{x}}=\operatorname{Ref}_{\mathbf{x}}$, which implies the equality $r_{\mathbf{x}}=\frac{1}{2}(\rho_{\mathbf{x}}+\operatorname{Pal}_{\mathbf{x}})$;
  • (b)
    or else there exists $n_{0}$ such that $\rho_{\mathbf{x}}(n)=\operatorname{Unr}_{\mathbf{x}}(n)$ for all $n\geq n_{0}$, which implies $r_{\mathbf{x}}(n)=\rho_{\mathbf{x}}(n)$ for all $n\geq n_{0}$.
Proof.
If $\mathbf{x}$ is reversal-closed, then $r_{\mathbf{x}}=\frac{1}{2}(\rho_{\mathbf{x}}+\operatorname{Pal}_{\mathbf{x}})$ from Theorem 9(b) above. Now suppose that $\mathbf{x}$ has an unreflected factor $w$. Since $\mathbf{x}$ is uniformly recurrent, every sufficiently long factor of $\mathbf{x}$ contains $w$ as a factor, which implies that this long factor itself is unreflected. This exactly says that $\operatorname{Ref}_{\mathbf{x}}(n)=0$ for $n$ large enough (and in particular $\operatorname{Pal}_{\mathbf{x}}(n)=0$ for $n$ large enough). This implies from Equalities (5) that, for $n$ large enough, $\rho_{\mathbf{x}}(n)=\operatorname{Unr}_{\mathbf{x}}(n)$, and so $r_{\mathbf{x}}(n)=\rho_{\mathbf{x}}(n)$. ∎
Now we exhibit sequences with particular behaviors of their reflection complexity.
Example 12.
It is possible to construct an aperiodic automatic sequence $\mathbf{x}$ such that $r_{\mathbf{x}}(n)=\rho_{\mathbf{x}}(n)$ and $\operatorname{Pal}_{\mathbf{x}}(n)>0$ for all $n$. An example of such a sequence is given by a fixed point of the morphism $0\mapsto 01$, $1\mapsto 23$, $2\mapsto 45$, $3\mapsto 23$, $4\mapsto 4$ 44, and $5\mapsto 44$. This sequence has no reflected factors except palindromes, and there is exactly one palindrome of each length $>1$.
Example 13.
Consider the sequence $\mathbf{x}$ on $\{0,1\}$ whose $n$th prefix $x_{n}$ is given recursively as follows: $x_{0}=01$ and $x_{n+1}=x_{n}01x_{n}^{R}$ for all $n\geq 0$. See [19, Section 3] or [15, Example 3.1]. The sequence $\mathbf{x}$ is uniformly recurrent, reversal-closed, $2$-automatic, and accepted by a DFAO of 6 states (see, e.g., [6]), and contains only a finite number of palindromes. Furthermore, for all sufficiently large $n$, we have $r_{\mathbf{x}}(n)=\dfrac{1}{2}\rho_{\mathbf{x}}(n)$.
Example 14.
It is also possible to construct an aperiodic automatic sequence where the only palindromes are of length 1, but there are reflected factors of each length $>1$. Let $g_{n}$ be the prefix of length $22^{n}-$ 2 of $(012)^{\omega}$. Then an example of an automatic sequence satisfying the desired properties is
[…]
where [80, Theorem 1] is required (observe that we intertwine the sequences $(3456)^{\omega}$ and $g_{1}g_{2}^{R}g_{3}g_{4}^{R}\cdots$ to build $\mathbf{x}$).
Example 15.
There is an automatic sequence $\mathbf{x}$ over the alphabet $\{0,1\}$ such that $\text{Ref}_{\mathbf{x}}(n)=\Omega(n)$ and such that $\text{Unr}_{\mathbf{x}}(n)=\Omega(n)$. Namely, consider the image under the coding $0,1,2\mapsto 0$ and $3,4\mapsto 1$ of the fixed point, starting with 00, of the morphism $0\mapsto 01$, $1\mapsto 23$, $2\mapsto 32$, $3\mapsto 42$, and $4\mapsto 43$.
Example 16.
We also provide a construction of an automatic sequence $\mathbf{x}$ such that $r_{\mathbf{x}}(n+1)<r_{\mathbf{x}}(n)$ for all odd $n\geq 3$. In particular, let $\mathbf{x}$ denote the sequence given by applying the coding $a,b,d\mapsto 1$ and $c\mapsto 0$ to the fixed point, starting with $a$, of the morphism defined by $a\mapsto ab$, $b\mapsto cd$, $c\mapsto cd$, and $d\mapsto bb$. This gives us sequence [92, A039982] in the OEIS. Computing the reflection complexity of $\mathbf{x}$ (e.g., using Walnut) gives that
$r_{\mathbf{x}}(n)=\begin{cases}n+1,&\text{for odd $n\geq 1$};\\ n-1,&\text{for even $n\geq 4$}.\end{cases}$
Actually we even have that $r_{\mathbf{x}}(n+1)=r_{\mathbf{x}}(n)-1$ for all odd $n\geq 3$. See Theorem 17.
With extra hypotheses on a sequence $\mathbf{x}$, we can give more precise results in comparing the respective growths of reflection and factor complexities. We will need Theorem 17 below. Note that Part (b) of this theorem was originally stated for uniformly recurrent sequences: see [14, Theorem 1.2]. However, its proof only requires the sequences to be recurrent (see [15, p. 449] and also [24, Footnote, p. 493]). Furthermore we have seen that a reversal-closed sequence must be recurrent (see Remark 10). Thus we can state the theorem as follows (also see Theorem 11).
Theorem 17.
  • (a)
    Let $\mathbf{x}$ be a uniformly recurrent sequence. If $\mathbf{x}$ is not closed under reversal, then $\operatorname{Pal}(n)=0$ for $n$ large enough (actually one even has $\operatorname{Ref}_{\mathbf{x}}(n)=0$ for $n$ large enough).
  • (b)
    Let $\mathbf{x}$ be a reversal-closed sequence. For all $n\geq 0$, we have
    $\operatorname{Pal}_{\mathbf{x}}(n+1)+\operatorname{Pal}_{\mathbf{x}}(n)\leq\rho_{\mathbf{x}}(n+1)-\rho_{\mathbf{x}}(n)+2.$
Remark 18.
There exist sequences that are uniformly recurrent, reversal-closed, and have no long palindromes (see [19]; also see Example 13 above).
We deduce the following results from Theorem 17.
Theorem 19.
Let $\mathbf{x}$ be a reversal-closed sequence. For all $n\geq 0$, we have
$\frac{1}{2}\rho_{\mathbf{x}}(n)\leq r_{\mathbf{x}}(n)\leq\frac{1}{2}\rho_{\mathbf{x}}(n+1)+1.$
Proof.
Using the first inequality in Theorem 9, the statement in Theorem 9(b), and Theorem 17(b), we have
$\frac{1}{2}\rho_{\mathbf{x}}(n)\leq r_{\mathbf{x}}(n)=\frac{1}{2}(\rho_{\mathbf{x}}(n)+\operatorname{Pal}_{\mathbf{x}}(n))\leq\frac{1}{2}(\rho_{\mathbf{x}}(n)+\rho_{\mathbf{x}}(n+1)-\rho_{\mathbf{x}}(n)+2)$
for all $n\geq 0$. The desired inequalities follow. ∎
Proposition 20.
Let $\mathbf{x}$ be a reversal-closed sequence. Then we have $r_{\mathbf{x}}(n+1)+r_{\mathbf{x}}(n)\leq\rho_{\mathbf{x}}(n+1)+1$ for all $n\geq 0$.
Proof.
It is enough to combine Theorems 17(b) and 9(b). ∎
On the other hand, we can use a result of [4] to obtain the following theorem.
Theorem 21.
Let $\mathbf{x}$ be a non-eventually periodic and reversal-closed sequence. For all $n\geq 1$, we have
$\frac{1}{2}\rho_{\mathbf{x}}(n)\leq r_{\mathbf{x}}(n)<\frac{1}{2}\rho_{\mathbf{x}}(n)+\frac{8}{n}\rho_{\mathbf{x}}\left(n+\left\lfloor\frac{n}{4}\right\rfloor\right).$
Proof.
Given a non-eventually periodic sequence $\mathbf{x}$, we have from [4, Theorem 12] the inequality
$\operatorname{Pal}_{\mathbf{x}}(n)<\frac{16}{n}\rho_{\mathbf{x}}\left(n+\left\lfloor\frac{n}{4}\right\rfloor\right)$
for all $n\geq 1$. The statement follows from this inequality and Theorem 9. ∎
Corollary 22.
Let $\mathbf{x}$ be a non-eventually periodic and reversal-closed sequence. If its factor complexity satisfies $\rho_{\mathbf{x}}(n+1)\sim\rho_{\mathbf{x}}(n)$ or $\frac{\rho_{\mathbf{x}}(2n)}{\rho_{\mathbf{x}}(n)}=o(n)$, then
$r_{\mathbf{x}}(n)\sim\frac{1}{2}\rho_{\mathbf{x}}(n)$
when $n$ tends to infinity. In particular, this equivalence holds if $\mathbf{x}$ is non-eventually periodic, reversal-closed, and morphic.
Proof.
Let $\mathbf{x}$ be a non-eventually periodic and reversal-closed sequence. If $\rho_{\mathbf{x}}(n+1)\sim\rho_{\mathbf{x}}(n)$, then, from Theorem 19, we obtain that $r_{\mathbf{x}}(n)\sim\frac{1}{2}\rho_{\mathbf{x}}(n)$ when $n$ tends to infinity. Now, if $\frac{\rho_{\mathbf{x}}(2n)}{\rho_{\mathbf{x}}(n)}=o(n)$, we obtain, from Theorem 21, and using the fact that $\rho_{\mathbf{x}}$ is non-decreasing, that
$\frac{1}{2}\rho_{\mathbf{x}}(n)\leq r_{\mathbf{x}}(n)<\frac{1}{2}\rho_{\mathbf{x}}(n)+\frac{8}{n}\rho_{\mathbf{x}}(2n)=\frac{1}{2}\rho_{\mathbf{x}}(n)+o(\rho_{\mathbf{x}}(n)),$
which is enough.
Now suppose that, in addition, the sequence $\mathbf{x}$ is morphic. We know that either $\rho_{\mathbf{x}}(n)=\Theta(n^{2})$ or $\rho_{\mathbf{x}}(n)=O(n^{3/2})$ (see [45, 44]). In the first case, then $\frac{\rho_{\mathbf{x}}(2n)}{\rho_{\mathbf{x}}(n)}$ is bounded, and hence $o(n)$. If $\rho_{\mathbf{x}}(n)=O(n^{3/2})$, since ${\mathbf{x}}$ is not eventually periodic (hence $\rho_{\mathbf{x}}(n)\geq n+1$), there exists a constant $C>0$ such that
$\frac{\rho_{\mathbf{x}}(2n)}{\rho_{\mathbf{x}}(n)}\leq C\frac{n^{3/2}}{n+1}=o(n).$
This finishes the proof. ∎
The upper bound in Theorem 21 raises questions about the growth properties of the function $r_{\mathbf{x}}$ more generally, apart from the case where the set of factors of $\mathbf{x}$ satisfies the hypotheses of Theorem 21. This leads us to the growth property in Theorem 23 below.
Theorem 23.
Let $\mathbf{x}$ be a sequence. Then $r_{\mathbf{x}}(n)\leq r_{\mathbf{x}}(n+2)$ for all $n\geq 0$.
Proof.
The result is clear for $n=0$, so assume $n>0$ in what follows. Let $c$ be a letter not in the alphabet of $\mathbf{x}$, and define ${\mathbf{y}}=c{\mathbf{x}}$. Then $r_{\mathbf{y}}(n)=r_{\mathbf{x}}(n)+1$ for all $n>0$, since $\mathbf{y}$ has exactly one additional factor for each length $n\geq 1$; namely, the prefix of length $n$. Thus, it suffices to prove the claim for $\mathbf{y}$ instead of $\mathbf{x}$.
With each length-$n$ factor $w$ of $\mathbf{y}$ associate a set $S_{w}$ of length-$(n+2)$ factors of $\mathbf{y}$, as follows: if $w$ is the length-$n$ prefix of $\mathbf{y}$, then $S_{w}:=\{w^{\prime}\}$, where $w^{\prime}$ is the prefix of length $n+2$ of $\mathbf{y}$. We call such a factor exceptional. Otherwise, define $S_{w}:=\{z\in\operatorname{Fac}({\mathbf{y}})\,:\,z=awb\text{ for some letters$a,b$}\}$. Note that the sets $S_{w}$, over all length-$n$ factors of $\mathbf{y}$, are pairwise disjoint, and cover all the length-$(n+2)$ factors of $\mathbf{y}$.
For a factor $w$ of $\mathbf{y}$, define $[w]_{1}=1$ if $w$ is a palindrome, and 00 otherwise. Similarly, $[w]_{2}=1$ if $w^{R}$ is not a factor of $\mathbf{y}$ and 00 otherwise. Finally, define $[w]_{3}=1$ if $w^{R}$ is also a factor of $\mathbf{y}$ but $w$ is not a palindrome, and 00 otherwise. Notice that these three cases are disjoint and subsume all possibilities for factors of $\mathbf{y}$ (also recall the decomposition at the beginning of the section). We extend this notation to sets by defining $[S]_{i}=\sum_{w\in S}[w]_{i}$ for $i\in\{1,2,3\}$. Define $[w]=[w]_{1}+[w]_{2}+[w]_{3}/2$ and similarly for $[S]$. From Equalities (5), we know that
$r_{\mathbf{y}}(n)=\sum_{\begin{subarray}{c}|w|=n\\ w\in\operatorname{Fac}({\mathbf{y}})\end{subarray}}[w]$
while
$r_{\mathbf{y}}(n+2)=\sum_{\begin{subarray}{c}|w|=n\\ w\in\operatorname{Fac}({\mathbf{y}})\end{subarray}}[S_{w}].$
Therefore, to show the desired inequality $r_{\mathbf{y}}(n)\leq r_{\mathbf{y}}(n+2)$, it suffices to show that $[w]\leq[S_{w}]$ for all length-$n$ factors $w$ of $\mathbf{y}$.
Suppose $w$ is exceptional. Recall that $w$ starts with $c$, which appears nowhere else in $\mathbf{y}$. Then $[w]_{1}=[w]_{3}=0$, but $[w]_{2}=1$. And $S_{w}=\{wab\}$, so $[S_{w}]_{1}=[S_{w}]_{3}=0$, but $[S_{w}]_{2}=1$. Therefore $[w]\leq[S_{w}]$.
Now suppose $w$ is not exceptional. There are three cases to consider, noting that when $n=1$, we have that $[w]_{1}=1$ and $[w]_{2}=0$.
Case 1: If $[w]_{1}=1$, then $w$ is a palindrome. Consider a factor $awb\in S_{w}$. If it is a palindrome, then $[awb]_{1}=1$, so $[w]\leq[awb]$. If $awb$ is not a palindrome, then $awb\not=(awb)^{R}=bw^{R}a=bwa$. Thus $a\not=b$. If $bwa$ is not a factor of $\mathbf{y}$, then $[awb]_{2}=1$, so $[w]\leq[awb]$. If $bwa$ is a factor of $\mathbf{y}$, then $bwa\in S_{w}$ and $[awb]_{3}+[bwa]_{3}=2$, so in all cases $[w]\leq[awb]$. Thus $[w]\leq[S_{w}]$.
Case 2: If $[w]_{2}=1$, then $w^{R}$ is not a factor of $\mathbf{y}$. Consider a factor $awb\in S_{w}$. Then $(awb)^{R}=bw^{R}a$, so $(awb)^{R}$ cannot be a factor of $y$ either. Hence $[awb]_{2}=1$, $[w]\leq[awb]$, and hence $[w]\leq[S_{w}]$.
Case 3: If $[w]_{3}=1$, then $w^{R}$ is a factor of $\mathbf{y}$, but $w$ is not a palindrome. Consider a factor $awb\in S_{w}$. If $awb$ is a palindrome, then $awb=(awb)^{R}=bw^{R}a$, so $w^{R}$ would be a palindrome, a contradiction. So $awb$ is not a palindrome and $[awb]_{1}=0$. If $(awb)^{R}=bw^{R}a$ is a factor of $\mathbf{y}$, then $[awb]_{3}=1$, so $[w]\leq[awb]$. If $(awb)^{R}$ is not a factor of $\mathbf{y}$, then $[awb]_{2}=1$, so $[w]<[awb]$. Thus $[w]\leq[S_{w}]$.
This completes the proof. ∎
Remark 24.
Another formulation of Theorem 23 above is that the sequence $(r_{\mathbf{x}}(n+1)+r_{\mathbf{x}}(n))_{n\geq 0}$ is non-decreasing.
With regard to the conjectures listed at the end of this section, we can prove a weaker form of Conjecture 28 for reversal-closed sequences, and a weaker form of Conjecture 27 for sequences without long palindromes. Also we can prove that Conjecture 29 holds for primitive morphic sequence.
Theorem 25.
Let $n_{0}\geq 0$ be an integer and let $\mathbf{x}$ be a sequence with no palindrome of length $\geq n_{0}$. Then $(r_{\mathbf{x}}(n))_{n\geq 0}$ is eventually non-decreasing: $r_{\mathbf{x}}(n)\leq r_{\mathbf{x}}(n+1)$ for $n\geq n_{0}$. Furthermore, if $r_{\mathbf{x}}(n+2)=r_{\mathbf{x}}(n)$ for some $n\geq n_{0}$, then the sequence $\mathbf{x}$ is eventually periodic.
Proof.
By combining the assumption and the second equality of Lemma 8, we have that
$r_{\mathbf{x}}(n)=\dfrac{1}{2}(\rho_{\mathbf{x}}(n)+\operatorname{Unr}_{\mathbf{x}}(n))$ (6)
for $n\geq n_{0}$. Since both $(\rho_{\mathbf{x}}(n))_{n\geq 0}$ and $(\operatorname{Unr}_{\mathbf{x}}(n))_{n\geq 0}$ are non-decreasing, we see that $(r_{\mathbf{x}}(n))_{n\geq n_{0}}$ is non-decreasing, which gives that $r_{\mathbf{x}}(n)\leq r_{\mathbf{x}}(n+1)$ for $n\geq n_{0}$. This shows the first part of the statement. For the second part, if we have $r_{\mathbf{x}}(n+2)=r_{\mathbf{x}}(n)$ for some $n\geq n_{0}$, then Equality (6) implies that $\rho_{\mathbf{x}}(n+2)+\operatorname{Unr}_{\mathbf{x}}(n+2)=\rho_{\mathbf{x}}(n)+\operatorname{Unr}_{\mathbf{x}}(n)$. Hence
$\rho_{\mathbf{x}}(n+2)+\operatorname{Unr}_{\mathbf{x}}(n+2)=\rho_{\mathbf{x}}(n+1)+\operatorname{Unr}_{\mathbf{x}}(n+1)=\rho_{\mathbf{x}}(n)+\operatorname{Unr}_{\mathbf{x}}(n).$ (7)
Hence $\rho_{\mathbf{x}}(n+1)=\rho_{\mathbf{x}}(n)$, which implies that $\mathbf{x}$ is eventually periodic from Theorem 5. ∎
Numerical experiments concerning the growth of the reflection complexity have led us to formulate Conjectures 2729 below. Actually, Theorems 11 and 25 can be used to show that Conjecture 28 holds if $\mathbf{x}$ is uniformly recurrent (also see Theorem 30 below). In the same vein, Theorem 11 and Corollary 22 imply the following corollary.
Corollary 26.
Conjecture 29 holds for non-eventually periodic primitive morphic sequences.
Proof.
Let $\mathbf{x}$ be a primitive morphic sequence. We know that $\mathbf{x}$ is uniformly recurrent. Thus, from Theorem 11, either $\mathbf{x}$ is reversal-closed, or else it has no long palindromes. If $\mathbf{x}$ is reversal-closed, then by Corollary 22, we have $r_{\mathbf{x}}(n)\sim\frac{1}{2}\rho_{\mathbf{x}}(n)$. Otherwise, $\mathbf{x}$ has no long palindromes, then, still from Theorem 11, we have that $r_{\mathbf{x}}(n)=\rho_{\mathbf{x}}(n)$ for $n$ large enough. ∎
We conclude this section by leaving the following conjectures as open problems.
Conjecture 27.
Let $\mathbf{x}$ be a sequence. Then $r_{\mathbf{x}}(n)=r_{\mathbf{x}}(n+2)$ for some $n$ if and only if $\mathbf{x}$ is eventually periodic.
Note that one direction is true. We have even more: namely, if the sequence $\mathbf{x}$ is eventually periodic, then $r_{\mathbf{x}}(n)=r_{\mathbf{x}}(n+2)$, for all $n$ large enough (see Theorem 39 below).
Conjecture 28.
Let $\mathbf{x}$ be a sequence of at most linear factor complexity. Then $r_{\mathbf{x}}(n+1)-r_{\mathbf{x}}(n)$ is bounded for all $n\geq 0$. Hence, in particular, if $\mathbf{x}$ is (generalized) automatic, so is $(r_{\mathbf{x}}(n+1)-r_{\mathbf{x}}(n))_{n\geq 0}$.
It can be shown that Conjecture 28 holds for the Thue–Morse, period-doubling, Golay–Shapiro, second-bit, paperfolding, Stewart choral, Baum-Sweet, Chacon, and Mephisto-Waltz sequences.
Conjecture 29.
Let $\mathbf{x}$ be a sequence. If the limit $\lim_{n\to\infty}\frac{r_{\mathbf{x}}(n)}{\rho_{\mathbf{x}}(n)}$ exists, then it is either equal to $\frac{1}{2}$ or to $1$.
4 Flye Sainte-Marie–Rauzy graphs for reflec-tion-equivalent classes
In this section, we prove the following inequality:
Theorem 30.
Let $\mathbf{x}$ be a sequence. Then $r_{\mathbf{x}}(n)\leq r_{\mathbf{x}}(n+1)+1$ for all $n\geq 0$.
Note that we can have equality in Theorem 30 for infinitely many values of $n$—see Example 16 above.
Initially, we discovered this growth property empirically. We prove it here by means of a graph construction related to the work of Flye Sainte-Marie. We remark that another recent application of this idea leads to a bound on the maximal number of powers in finite words [23, 64].
Flye Sainte-Marie graphs (also known as Rauzy graphs or De Bruijn graphs) for sequences $\mathbf{x}$ can be defined in the following way: for an integer $n\geq 0$, the graph $\Gamma_{\mathbf{x}}(n)$ is a directed graph whose vertex set is $\operatorname{Fac}_{\mathbf{x}}(n)$ and edge set is $\operatorname{Fac}_{\mathbf{x}}(n+1)$. An edge $e\in\operatorname{Fac}_{\mathbf{x}}(n+1)$ starts at the vertex $u$ and ends at the vertex $v$ if and only if $u$ is a prefix of $e$ and $v$ is a suffix. To simplify the notation, for words $w_{1},w_{2}$ of length $n$ and a word $w_{3}$ of length $n+1$, we write $w_{1}\xrightarrow{w_{3}}w_{2}$ if $w_{1}$ and $w_{2}$ are a prefix and a suffix of $w_{3}$, respectively.
Remark 31.
Flye Sainte-Marie first introduced the previous graph construction in [54] to represent the overlaps between all the binary words of the same length. De Bruijn rediscovered the same graph construction in [25] and acknowledged Flye Sainte-Marie’s priority of [54] later in [26]. Rauzy used the same construction in [81] to study the overlaps of factors in a word.
A directed graph is (weakly) connected if there is an undirected path between any pair of its vertices.
The following proposition is “folklore”. To the authors’ knowledge, the proof of this proposition does not seem to properly appear in the literature.
Proposition 32.
Let $\mathbf{x}$ be a sequence. For all $n\geq 0$, the graph $\Gamma_{\mathbf{x}}(n)$ is connected.
Proof.
Let $u$ and $v$ be two distinct length-$n$ factors of $\mathbf{x}$ and let us suppose that $u$ occurs before $v$ in $\mathbf{x}$. Then there are two integers $i<j$ such that $\mathbf{x}[i..i+n-1]=u$ and $\mathbf{x}[j..j+n-1]=v$. Then the path $\mathbf{x}[i..i+n],\mathbf{x}[i+1..i+n+1],\ldots,\mathbf{x}[j-1..j+n-1]$ connects $u$ and $v$. ∎
Remark 33.
Proposition 32 has already been (indirectly) mentioned in [85, p. 199, lines 4–6], which states that every graph $\Gamma_{\mathbf{x}}(n)$ necessarily has a vertex $o$ from which every other vertex can be reached by a direct path. Thus each pair of vertices $u$, $v$ can be connected by an undirected path passing through $o$.
Let $\mathbf{x}$ be a sequence. For all $n\geq 0$, let $\operatorname{ClassRef}_{\mathbf{x}}(n)$ be the set of all reflection-equivalent classes of length-$n$ factors of $\mathbf{x}$, i.e.,
$\operatorname{ClassRef}_{\mathbf{x}}(n)=\{\{u,u^{R}\}\,:\,u\in\operatorname{Fac}_{\mathbf{x}}(n)\},$
where $u$ may be equal to $u^{R}$. From the definition, the cardinality of the set $\operatorname{ClassRef}_{\mathbf{x}}(n)$ is $r_{\mathbf{x}}(n)$.
[…]
Remark 34.
From the definition, each element in $\operatorname{Fac}_{\mathbf{x}}(n+1)$ appears exactly once as an edge in $\Lambda_{\mathbf{x}}(n)$. This is easily seen by letting $e\in\operatorname{Fac}_{\mathbf{x}}(n+1)$ and by letting $u$ and $v$ be the length-$n$ prefix and suffix of $e$, respectively, and by using the property that $\{u,u^{R}\},\{v,v^{R}\}\in\operatorname{ClassRef}_{\mathbf{x}}(n)$. Thus, the number of vertices in $\Lambda_{\mathbf{x}}(n)$ is $r_{\mathbf{x}}(n)$ and the number of edges in $\Lambda_{\mathbf{x}}(n)$ is $\rho_{\mathbf{x}}(n+1)$.
Remark 35.
Letting $\mathbf{x}$ be a sequence, the graph $\Lambda_{\mathbf{x}}(n)$ is connected for arbitrary $n\geq 0$. This follows in a direct way from $\Gamma_{\mathbf{x}}(n)$ being connected from Proposition 32.
Lemma 36.
Let $\mathbf{x}$ be a sequence. Let $u,v$ be two distinct length-$(n+1)$ factors of $\mathbf{x}$ for some non-negative integer $n$. Then $u=v^{R}$ if and only if there are two distinct vertices $c_{1},c_{2}$ in $\Lambda_{\mathbf{x}}(n)$ such that $u$ is an edge from $c_{1}$ to $c_{2}$ and $v$ is an edge from $c_{2}$ to $c_{1}$.
Remark 37.
If there are two distinct elements in $c_{1}$ and/or $c_{2}$, there may be at most two distinct edges from $c_{1}$ to $c_{2}$ and vice-versa. So, to prove Lemma 36, we need to prove that for two distinct vertices, there is at most one edge from one to the other. Furthermore, if these two vertices are connected in both directions, then the two connecting edges in between are reflectively equivalent. Graphically, Lemma 36 gives us that $u=v^{R}$ and $u\neq v$ if and only if the graph $\Lambda_{\mathbf{x}}(n)$ has an occurrence of the pattern depicted in Figure 1.
$c_{1}$$c_{2}$$v$$u$
Figure 1: The consequence of Lemma 36 in the graph $\Lambda_{\mathbf{x}}(n)$, when $u=v^{R}$ but $u\not=v$.
Proof of Lemma 36.
If $u$ and $v=u^{R}$ are both edges of $\Lambda_{\text{{\bf x}}}(n)$, then we now show that $u=(c_{1},c_{2})$ and $v=(c_{2},c_{1})$ for some vertices $c_{1}$, $c_{2}$ of $\Lambda_{\text{{\bf x}}}(n)$ (see Fig. 1).
As indicated in Remark 37, we need to show that, in the case where $c_{1}\neq c_{2}$, there are at most two edges between two distinct vertices, and if there are exactly two edges, then one is the reverse of the other. In this direction, the given claim is clearly true for $n=0$ and $n=1$. For $n\geq 2$, we now show that if there exists $u$ such that $u\neq u^{R}$ and $u,u^{R}\in\operatorname{Fac}_{\mathbf{x}}(n+1)$, then there exist two distinct vertices satisfying the conditions in the statement. Let $u=au^{\prime}b$, where $a,b$ are letters. From the hypothesis that $u^{R}\in\operatorname{Fac}_{\mathbf{x}}(n+1)$, one has $au^{\prime},u^{\prime}b,u^{\prime R}a,bu^{\prime R}\in\operatorname{Fac}_{\mathbf{x}}(n)$. Thus, $c_{1}=\{au^{\prime},u^{\prime R}a\}$ and $c_{2}=\{u^{\prime}b,bu^{\prime R}\}$ are two vertices in the graph $\Lambda_{\mathbf{x}}(n)$. Moreover, since
$au^{\prime}\xrightarrow{au^{\prime}b}u^{\prime}b\text{ and }bu^{\prime R}\xrightarrow{bu^{\prime R}a}u^{\prime R}a,$
the word $au^{\prime}b$ is an edge from $c_{1}$ to $c_{2}$ and $bu^{\prime R}a$ is an edge from $c_{2}$ to $c_{1}$. Finally, to prove that $c_{1}\neq c_{2}$, we can use a proof by contradiction together with a case analysis by considering the $au^{\prime}=bu^{\prime R}$ case and the $au^{\prime}=u^{\prime}b$ case. A similar approach can be used to prove that $u=v^{R}$. We leave the details to the reader. ∎
Now we introduce the Flye Sainte-Marie graphs for reflection-equivalent classes. Let $\mathbf{x}$ be a sequence. For all $n\geq 0$, let $K_{\mathbf{x}}(n)$ be a directed graph such that its vertex set is $\operatorname{ClassRef}_{\mathbf{x}}(n)$ and its edge set is $\operatorname{ClassRef}_{\mathbf{x}}(n+1)$. Let $e\in\operatorname{ClassRef}_{\mathbf{x}}(n+1)$. We define the edges of $K_{\mathbf{x}}(n)$ as follows:
  • if $|e\cap\operatorname{Fac}_{\mathbf{x}}(n+1)|=1$, there exist a unique word $u\in e$ and two vertices $c_{1},c_{2}\in\operatorname{ClassRef}_{\mathbf{x}}(n)$ such that $u$ is from $c_{1}$ to $c_{2}$ in $\Lambda_{\mathbf{x}}(n)$. In the graph $K_{\mathbf{x}}(n)$, let $e$ be the edge from $c_{1}$ to $c_{2}$.
  • if $|e\cap\operatorname{Fac}_{\mathbf{x}}(n+1)|=2$, from Lemma 36 there exists a unique word $u\in e$ satisfying the following four conditions:
    • $u\neq u^{R}$;
    • $u$ is lexicographically smaller than $u^{R}$;
    • $u,u^{R}\in\operatorname{Fac}_{\mathbf{x}}(n+1)$;
    • there exist two vertices $c_{1},c_{2}\in\operatorname{ClassRef}_{\mathbf{x}}(n)$ such that $u$ is from $c_{1}$ to $c_{2}$ and $u^{R}$ is from $c_{2}$ to $c_{1}$ in $\Lambda_{\mathbf{x}}(n)$.
    In the graph $K_{\mathbf{x}}(n)$, let $e$ be the edge from $c_{1}$ to $c_{2}$.
Proposition 38.
Let $\mathbf{x}$ be a sequence. For all $n\geq 0$, the graph $K_{\mathbf{x}}(n)$ has exactly $r_{\mathbf{x}}(n)$ vertices and $r_{\mathbf{x}}(n+1)$ edges. Moreover, it is connected.
Proof.
From the construction, each element in $\operatorname{ClassRef}_{\mathbf{x}}(n+1)$ appears exactly once as an edge in $K_{\mathbf{x}}(n)$ and the vertex set of this graph is exactly $\operatorname{ClassRef}_{\mathbf{x}}(n)$. Thus, the number of vertices and edges is respectively $r_{\mathbf{x}}(n)$ and $r_{\mathbf{x}}(n+1)$. For the connectivity of the graph $K_{\mathbf{x}}(n)$, it is enough to show that every pair of vertices $c_{1},c_{2}$ connected by a single edge $u$ in $\Lambda_{\mathbf{x}}(n)$ is also connected by the single edge $\{u,u^{R}\}$ in $K_{\mathbf{x}}(n)$. Since the graph $\Lambda_{\mathbf{x}}(n)$ is connected, the graph $K_{\mathbf{x}}(n)$ is also connected. ∎
We can now complete the proof of Theorem 30:
Proof of Theorem 30.
This follows immediately from Proposition 38, since $K_{\mathbf{x}}(n)$ is a connected graph with $r_{\mathbf{x}}(n)$ vertices and $r_{\mathbf{x}}(n+1)$ edges. Thus $r_{\mathbf{x}}(n)\leq r_{\mathbf{x}}(n+1)+1$. ∎
5 Eventually periodic sequences
We can characterize eventually periodic sequences (i.e., sequences that are periodic from some index on) in terms of their reflection complexity.
Theorem 39.
A sequence $\mathbf{x}$ is eventually periodic if and only if both sequences $(r_{\mathbf{x}}(2n))_{n\geq 0}$ and $(r_{\mathbf{x}}(2n+1))_{n\geq 0}$ are eventually constant.
Proof.
From Theorem 23 both sequences $(r_{\mathbf{x}}(2n))_{n\geq 0}$ and $(r_{\mathbf{x}}(2n+1))_{n\geq 0}$ are non-decreasing. Also, from the inequalities $\frac{1}{2}\rho_{\mathbf{x}}(n)\leq r_{\mathbf{x}}(n)\leq\rho_{\mathbf{x}}(n)$ in Theorem 9, and the fact that the sequence $(\rho_{\mathbf{x}}(n))_{n\geq 0}$ is non-decreasing, we have that either the three integer sequences $(r_{\mathbf{x}}(2n))_{n\geq 0}$, $(r_{\mathbf{x}}(2n+1))_{n\geq 0}$, and $(\rho_{\mathbf{x}}(n))_{n\geq 0}$ are all bounded, or else none of them is. Furthermore, we know that $(\rho_{\mathbf{x}}(n))_{n\geq 0}$ is bounded if and only if the sequence $\mathbf{x}$ is eventually periodic (Theorem 5 above). Hence, we have two cases depending on the periodicity of $\mathbf{x}$.
  • (a)
    If $\mathbf{x}$ is eventually periodic, then $(\rho_{\mathbf{x}}(n))_{n\geq 0}$ is bounded, so $(r_{\mathbf{x}}(2n))_{n\geq 0}$ and $(r_{\mathbf{x}}(2n+1))_{n\geq 0}$ are eventually constant.
  • (b)
    If $\mathbf{x}$ is not eventually periodic, its factor complexity is not bounded, thus both sequences $(r_{\mathbf{x}}(2n))_{n\geq 0}$ and $(r_{\mathbf{x}}(2n+1))_{n\geq 0}$ tend to infinity.
This ends the proof. ∎
Remark 40.
If $\mathbf{x}$ is eventually periodic, the eventual values of $(r_{\mathbf{x}}(2n))_{n\geq 0}$ and $(r_{\mathbf{x}}(2n+1))_{n\geq 0}$ can be either equal or distinct, as seen from the examples of the sequences $(01)^{\omega}$ and $(011)^{\omega}$.
6 Sturmian sequences and generalizations
In this section, we study Sturmian sequences as well as some generalizations. First we state the following result, which characterizes Sturmian sequences in terms of their reflection complexity.
Theorem 41.
Let $\mathbf{x}$ be a non-eventually periodic sequence over a finite alphabet.
  • (a)
    For all $n\geq 1$, we have $r_{\mathbf{x}}(n)\geq 1+\lfloor\frac{n+1}{2}\rfloor$;
  • (b)
    We have $r_{\mathbf{x}}(n)=1+\lfloor\frac{n+1}{2}\rfloor$ if and only if $\mathbf{x}$ is Sturmian.
Proof.
For each integer $n\geq 1$, let ${\mathcal{S}}_{n}$ be the permutation group on $n$ elements. Let $\sigma_{n}$ be the permutation defined by
$\sigma_{n}:=\begin{pmatrix}1&2&\ldots&n-1&n\\ n&n-1&\ldots&2&1\\ \end{pmatrix}$
and $G_{n}$ be the subgroup of ${\mathcal{S}}_{n}$ generated by $\sigma_{n}$, i.e., the group $\{\sigma_{n},{\rm id}_{n}\}$. The number of distinct orbits of $\{1,2,\ldots,n\}$ under $G_{n}$ is equal to $n/2$ if $n$ is even, and to $(n+1)/2$ if $n$ is odd, which can be written $\lfloor(n+1)/2\rfloor$ in both cases. Thus, applying [36, Theorem 1] proves the first item of the theorem and the implication $\Longrightarrow$ of the second item.
To prove the last assertion, suppose that $\mathbf{x}$ is a Sturmian sequence. We know that every Sturmian sequence is reversal-closed (see [47, Theorem 4, p. 77], where reversals are called mirror images). Furthermore, it is proved in [47, Theorem 5, p. 77] that a sequence is Sturmian if and only if it has one palindrome of all even lengths and two palindromes of all odd lengths. Now, from Theorem 9(b) we have that
$r_{\mathbf{x}}(n)=\frac{1}{2}(\rho_{\mathbf{x}}(n)+\operatorname{Pal}_{\mathbf{x}}(n))=\begin{cases}\frac{n+2}{2}=1+\lfloor\frac{n+1}{2}\rfloor,&\text{if $n$ even;}\\ \frac{n+3}{2}=1+\lfloor\frac{n+1}{2}\rfloor,&\text{if $n$ odd}.\end{cases}$
This completes the proof. ∎
With regard to the above referenced work of Charlier et al. [36], also see the related and recent work by Luchinin and Puzynina [67].
The following is an analog of the Morse–Hedlund theorem (which is recalled in Theorem 5 above).
Corollary 42.
A sequence $\mathbf{x}$ is eventually periodic if and only if there exists $n\geq 1$ such that $r_{\mathbf{x}}(n)\leq\left\lfloor\frac{n+1}{2}\right\rfloor$. Furthermore both sequences $(r_{\mathbf{x}}(2n))_{n}$ and $(r_{\mathbf{x}}(2n+1))_{n}$ are then eventually constant.
Proof.
Let $\mathbf{x}$ be a sequence. Contraposing Property (a) of Theorem 41, we obtain that if $r_{\mathbf{x}}(n)\leq\lfloor\frac{n+1}{2}\rfloor$ for some $n$, then $\mathbf{x}$ must be eventually periodic. Conversely, if $\mathbf{x}$ is eventually periodic, it has a bounded number of factors, hence there exists some integer $n$ for which the inequality of the statement is true. The last assertion is Theorem 39 above. ∎
Recall that every quasi-Sturmian sequence $\mathbf{x}$ can be written as $\mathbf{x}=yf(\mathbf{z})$, where $y$ is a word over a finite alphabet, $\mathbf{z}$ is a (necessarily binary) Sturmian sequence, and $f$ an aperiodic morphism from $\{0,1\}$ to a finite alphabet, see [31, 40, 78]. (Recall that $f$ aperiodic means that $f(01)\neq f(10)$.) We state the following theorem.
Theorem 43.
Let $\mathbf{x}=yf(\mathbf{z})$ be a quasi-Sturmian sequence, where $y$ is a word, $\mathbf{z}$ is a Sturmian sequence, and $f$ is an aperiodic morphism from $\{0,1\}$ to a finite alphabet. Then
  • (a)
    either $f(\mathbf{z})$ is reversal-closed and $r_{\mathbf{x}}(n)=\frac{n}{2}+O(1)$;
  • (b)
    or else $f(\mathbf{z})$ is not reversal-closed and $r_{\mathbf{x}}(n)=n+O(1)$.
Proof.
Since $r_{\mathbf{x}}(n)=r_{f(\mathbf{z})}(n)+O(1)$, it suffices to prove both statements for $r_{f(\mathbf{z})}$ instead of $r_{\mathbf{x}}$. Since $\mathbf{z}$ and, hence, $f(\mathbf{z})$ are both uniformly recurrent, we can apply Theorem 11 to $f(\mathbf{z})$. The desired result then easily follows by considering the case whereby $f(\mathbf{z})$ is reversal-closed, so that $r_{f(\mathbf{z})}=\frac{1}{2}(\rho_{f(\mathbf{z})}+\operatorname{Pal}_{f(\mathbf{z})})$, and the case such that it is not reversal-closed, so that $r_{f(\mathbf{z})}(n)=\rho_{f(\mathbf{z})}(n)$. ∎
Among several generalizations of Sturmian sequences, episturmian sequences have in particular the property—sometimes even taken as part of their definition—to be reversal-closed. Furthermore, their palindrome complexity has been studied [18, 55]. Here we develop a theorem similar to Theorem 41 above for these sequences.
Definition 44.
Let $A$ be a finite alphabet with cardinality $\ell$. A sequence $\mathbf{x}$ over $A$ is episturmian if it is reversal-closed and has at most one left special factor of each length. An episturmian sequence $\mathbf{x}$ is $\ell$-strict if it has exactly one left special factor of each length and every left special factor $u$ of $\mathbf{x}$ has $\ell$ distinct left extensions in $\mathbf{x}$.
We compute the reflection complexity of episturmian sequences as follows. (Recall that the factor complexity of an $\ell$-strict episturmian sequence is given by $\rho_{\mathbf{x}}(n)=(\ell-1)n+1$.)
Theorem 45.
Let $\mathbf{x}$ be an $\ell$-strict episturmian sequence. Then, for all $n\geq 0$, we have
$r_{\mathbf{x}}(n)=(\ell-1)\left\lfloor\frac{n+1}{2}\right\rfloor+1.$
Proof.
Let $\mathbf{x}$ be an $\ell$-strict episturmian sequence. The case $n=0$ is true. Assume that $n\geq 1$. Then by [46, Theorem 7], we have $\rho_{\mathbf{x}}(n)=(\ell-1)n+1$. We also know from [47] that
$\operatorname{Pal}_{\mathbf{x}}(n)=\begin{cases}1,&\text{if $n$ is even};\\ \ell,&\text{if $n$ is odd}.\end{cases}$
Using these facts, together with Theorem 9(b), we deduce the desired result. ∎
Example 46.
For the Tribonacci sequence $\mathbf{tr}$, which is the fixed point of the morphism $0\mapsto 01,1\mapsto 02,2\mapsto 0$, we have $r_{\mathbf{tr}}(n)=2\left\lfloor\frac{n+1}{2}\right\rfloor+1$ for all $n\geq 0$.
Since one interpretation of Sturmian sequences is the binary coding of irrational trajectories on a square billiard table, we turn our attention toward irrational trajectories on a hypercube. Using results due to Baryshnikov [16] together with Corollary 22, it follows in a direct way that
$r_{\mathbf{x}}(n)\sim\frac{1}{2}\sum_{k=0}^{\min{(s,n)}}k!\binom{s}{k}\binom{n}{k},$
when $n$ tends to infinity, for an irrational billiard sequence $\mathbf{x}$ on a hypercube of dimension $(s+1)$. In particular, we have that $r_{\mathbf{x}}(n)=\Theta(n^{s})$ when $n$ tends to infinity.
So-called complementation-symmetric Rote sequences, which were defined and studied by Rote in 1994 [85], are related to Sturmian sequences as stated below in Theorem 48.
Definition 47.
Let $\mathbf{x}$ be a binary sequence. Then $\mathbf{x}$ is called a Rote sequence if its factor complexity satisfies $\rho_{\mathbf{x}}(n)=2n$ for all $n\geq 1$. The sequence $\mathbf{x}$ is said to be complementation-symmetric if its set of factors is closed under the exchange morphism, i.e., if $w$ is a factor of $\mathbf{x}$, so is $E(w)$.
We consider the mapping $\Delta\colon\{0,1\}^{+}\to\{0,1\}^{*}$ defined as follows: $\Delta(a)=a$ for all $a\in\{0,1\}$ and
$\Delta(v(0)v(1)\cdots v(n))=u(0)u(1)\cdots u(n-1)$
for $n\geq 1$ with $u(i)=(v(i+1)-v(i))\bmod{2}$ for all $i\in\{0,\ldots,n-1\}$. There is a natural extension of $\Delta$ to sequences: if $\mathbf{x}=(x(n))_{n\geq 0}$ is a binary sequence, then $\Delta(\mathbf{x})$ is the sequence whose $n$th letter is defined by
$(x(n+1)-x(n))\bmod{2}$
for all $n\geq 0$. Observe that $\Delta(\mathbf{x})$ is the sequence of first differences of $\mathbf{x}$, taken modulo $2$.
Theorem 48 ([85]).
A binary sequence $\mathbf{x}$ is a complementation-symmetric Rote sequence if and only if $\Delta(\mathbf{x})$ is Sturmian.
In fact, for each Sturmian sequence $\mathbf{s}$, there are two associated complem-entation-symmetric Rote sequences $\mathbf{x}$ and $\mathbf{x}^{\prime}$ with $\mathbf{x}^{\prime}=E(\mathbf{x})$. The factors in $\mathbf{s}$ and its corresponding Rote sequences are closely related as shown below.
Proposition 49 ([85]; also see [68, Proposition 2] or [69, Lemma 2.7]).
Let $\mathbf{s}$ be a Sturmian sequence and let $\mathbf{x}$ be the complementation-symmetric Rote sequence such that $\mathbf{s}=\Delta(\mathbf{x})$. Then $u$ is a factor of $\mathbf{s}$ if and only if both words $v,v^{\prime}$ such that $u=\Delta(v)=\Delta(v^{\prime})$ are factors of $\mathbf{x}$. Furthermore, for every $n\geq 0$, $u$ occurs at position $n$ in $\mathbf{s}$ if and only if $v$ or $v^{\prime}$ occurs at position $n$ in $\mathbf{x}$.
Lemma 50.
A complementation-symmetric Rote sequence is reversal-closed.
Proof.
Let $\mathbf{x}$ be a complementation-symmetric Rote sequence. Let $\mathbf{s}$ be the Sturmian sequence corresponding to $\mathbf{x}$, i.e., $\mathbf{s}=\Delta(\mathbf{x})$ given by Theorem 48. Consider a factor $v$ of $\mathbf{x}$. Write $u=\Delta(v)$. Since $\mathbf{s}$ is reversal-closed, the word $u^{R}$ is also a factor of $\mathbf{s}$. Let $w$ and $w^{\prime}$ be the binary words such that $u^{R}=\Delta(w)=\Delta(w^{\prime})$ and $w^{\prime}=E(w)$. By Proposition 49, both $w$ and $w^{\prime}$ are factors of $\mathbf{x}$. Now observe that we have either $v^{R}=w$ or $v^{R}=w^{\prime}$. This ends the proof. ∎
We compute the reflection complexity of Rote sequences as follows.
Theorem 51.
Let $\mathbf{x}$ be a complementation-symmetric Rote sequence. Then its reflection complexity satisfies $r_{\mathbf{x}}(n)=n+1$ for all $n\geq 0$.
Proof.
Let $\mathbf{x}$ be a complementation-symmetric Rote sequence. We clearly have $r_{\mathbf{x}}(0)=1$. Now, for $n\geq 1$[4, Theorem 8] states that $\operatorname{Pal}_{\mathbf{x}}(n)=2$. We finish the proof using Lemma 50 and Theorem 9(b). ∎
7 Rich reversal-closed sequences
Rich sequences have several equivalent definitions. It is known that a word $w$ contains at most $|w|+1$ palindromic factors [46]. A sequence is called rich if each factor contains the maximal number of palindromic factors.
Theorem 52.
Let $\mathbf{x}$ be a reversal-closed sequence. Then $\mathbf{x}$ is rich if and only if $r_{\mathbf{x}}(n+1)+r_{\mathbf{x}}(n)=\rho_{\mathbf{x}}(n+1)+1$ for all $n\geq 0$.
Proof.
From [27, Theorem 1.1], the sequence $\mathbf{x}$ is rich if and only if the inequality in Theorem 17(b) is an equality. The result then follows from Theorem 9(b). ∎
Those among binary quasi-Sturmian sequences that are codings of rotations are rich, see [20, Theorem 19].
Corollary 53.
Let $\mathbf{x}$ be a binary reversal-closed quasi-Sturmian sequence. There exists a constant $C$ such that $r_{\mathbf{x}}(n+1)+r_{\mathbf{x}}(n)=n+C$ for $n$ large enough.
Proof.
Let $C^{\prime}$ be a constant such that $\rho_{\mathbf{x}}(n)=n+C^{\prime}$ for $n$ large enough. It is enough to choose $C=C^{\prime}+2$. ∎
Using Theorem 9(b) and [87, Corollaries 2.27 and 2.29], it is possible to bound the reflection complexity of rich sequences as follows.
Proposition 54.
Let $\mathbf{x}$ be a rich sequence over an alphabet of $q$ letters and write $\delta=\frac{2}{3(\log 3-\log 2)}$. Then $r_{\mathbf{x}}(n)\leq\frac{nq}{2}(2q^{2}n)^{\delta\log n}(1+nq^{3}(2q^{2}n)^{\delta\log n})$ for all $n\geq 1$.
Other sequences have a reflection complexity satisfying the equality of Theorem 52. For instance, it is the case for complementation-symmetric sequences, sequences canonically associated with some specific Parry numbers, and sequences coding particular interval exchange transformations. For more details, see [14, Section 3].
8 Automatic sequences
In this section, we study the reflection complexity of automatic sequences. First, in a positional numeration system $U$ having an adder (i.e., addition is recognizable by an automaton reading $U$-representations), we show that if a sequence is $U$-automatic, then its reflection complexity is a $U$-regular sequence. Furthermore we show how to effectively compute a linear representation for the sequence, making use of the free software Walnut [74, 90]. Next, we explore the reflection complexity of some famous automatic sequences, namely the Thue–Morse, the period-doubling, generalized paperfolding, generalized Golay–Shapiro, and the Baum–Sweet sequences.
8.1 Reflection complexity is computably regular
We now show that the reflection complexity of an automatic sequence is regular.
Theorem 55.
Let $U=(U(n))_{n\geq 0}$ be a positional numeration system such that there is an adder, and let $\mathbf{x}$ be a $U$-automatic sequence. Then $(r_{\mathbf{x}}(n))_{n\geq 0}$ is a $U$-regular sequence. Furthermore, a linear representation for $(r_{\mathbf{x}}(n))_{n\geq 0}$ is computable from the DFAO for $\mathbf{x}$.
Proof.
Here is a sketch of the proof before we give the details: We create a first-order logical formula asserting that the factor $\mathbf{x}[i..i+n-1]$ is the first occurrence of this factor, or its reversal. Then the number of such $i$ is precisely the reflection complexity at $n$. From this, we can create a linear representation for the number of such $i$.
Now some more details. We define the following logical formulas:
$\operatorname{{\textsc{FactorEq}}}(i,j,n)$ $:=\forall t\ (t<n)\implies\mathbf{x}[i+t]=\mathbf{x}[j+t]$
$\operatorname{{\textsc{FactorRevEq}}}(i,j,n)$ $:=\forall t\ (t<n)\implies\mathbf{x}[i+t]=\mathbf{x}[(j+n)-(t+1)]$ (8)
$\operatorname{{\textsc{RefComp}}}(i,n)$ $:=\forall j\ (j<i)\implies((\neg\operatorname{{\textsc{FactorEq}}}(i,j,n))$
($\ \wedge\ (\neg\operatorname{{\textsc{FactorRevEq}}}(i,j,n))).$
[…]
Once we have a linear representation for the reflection complexity, we can easily compute it for a given $n$. Furthermore, we can compare it to a guessed formula, provided that this formula can also be expressed as a linear representation (see [90]). In the next section we carry this out in detail for a number of famous sequences.
First we study the case of the Thue-Morse sequence. We want to emphasize the fact that the point of this example is not to reprove things that could be done more easily by appealing to existing theorems. It is to illustrate how the approach via Walnut can, in principle, carry out the various constructions for an arbitrary word, using Thue-Morse as an example simple enough where the matrices can actually be displayed.
We can compute a linear representation for the reflection complexity $r_{\mathbf{t}}(n)$ of the $2$-automatic Thue–Morse sequence $\mathbf{t}$, using the same approach as in the preceding section. Here we use the following Walnut code:
def factoreq_tm "At (t<n) => T[i+t]=T[j+t]"::
def factorreveq_tm "At (t<n) => T[i+t]=T[(j+n)-(t+1)]"::
def rc_tm n "Aj (j<i) => ((~\$factoreq_tm(i,j,n))
 & (~\$factorreveq_tm(i,j,n)))"::
This generates a linear representation of rank $66$, which can be minimized to the following.
$v=\left[\begin{array}[]{ccccccccc}1&0&0&0&0&0&0&0&0\end{array}\right],\quad w=\left[\begin{array}[]{ccccccccc}1&2&3&4&6&6&10&10&13\end{array}\right]^{T}$
μ(0)=133[
3300000000003300000000003300000000003300000000003300-26002310-103600-573306-665100-793333-5-28285100-7203318-181854
,  μ(1) = 133[ 033000000000033000000000033000000000033000-240039-661800-400043-10103000-7833333-36364200-8633335-38384800-720051-181854] .
Recall that Brlek [21], de Luca and Varricchio [42], and Avgustinovich [11] independently gave a simple recurrence for the number of length-n factors of t, namely =ρt(2n)+ρt(n)ρt(+n1) and =ρt(+2n1)2ρt(+n1) for n2. As it turns out, there is a simple relationship between rt and ρt.
Theorem 56.
Let t be the Thue–Morse sequence.
  • (a)
    For all n0, we have =rt(+2n1)ρt(+n1).
  • (b)
    For all n2, we have
    =rt(2n){+ρt(+n1)1,if m0 with +34-m11n4m;ρt(+n1),otherwise.
  • (c)
    There is an automaton of 14 states that computes the first difference (-rt(+n1)rt(n))n0.
Proof.
We prove each item separately.
  • (a)
    Above in Equalities (11) we computed a linear representation for rt(n). From this linear representation we can easily compute one for rt(+2n1) merely by replacing w with μ(1)w. (Indeed, base-2 representations of integers +2n1 all end with 1.)
    Next, we can compute a linear representation for ρt(+n1) using the following Walnut command.
    def sc_tm_offset n "Aj (j<i) => ~\$factoreq_tm(i,j,n+1)":
    This creates a linear representation of rank 6.
    Finally, we use a block matrix construction to compute a linear representation for the difference -rt(+2n1)ρt(+n1) and minimize it; the result is the 0 representation. This computation gives a rigorous proof of item (a).
  • (b)
    This identity can be proven in a similar way. We form the linear representation for
    rt(2n)-ρt(n+1)-[m:34-m1+1n4m],
    where the last term uses the Iverson bracket. We then minimize the result and obtain the 0 representation.
  • (c)
    We can compute a linear representation for the first difference (-rt(+n1)rt(n))n0, and then use the “semigroup trick” [90, Section 4.11] to prove that the difference is bounded and find the automaton for it. It is displayed in Figure 2.
    Figure 2: Automaton computing (-rt(+n1)rt(n))n0 where t is the Thue–Morse sequence.
These computations rigorously prove the three items of the claim. ∎
A similar approach can be used to evaluate the reflection complexity for the period-doubling sequence p, which may be defined as the fixed point of the morphism 001 and 100. This gives us that p is 2-automatic as well. By analogy with Theorem 56, we can show that for all n0, we have =rp(+2n1)+ρp(n)1, and, for all n2, we have
=rp(2n){-ρp(+n1)1,if m0 with 32-m1n-2+m11;-ρp(+n1)2,otherwise,
and we may similarly devise an analogue of part (c) of Theorem 56.
A paperfolding sequence pf is a binary sequence p1p2p3 specified by a sequence of binary unfolding instructions f0f1f2, as the limit of the sequences pf0f1f2, defined as follows:
pε=εandpf0f+i1=pf0fif+i1E(pf0fiR) for all i0
where E is the exchange morphism. For example, if =f000, we get the simplest paperfolding sequence
=p0010011000110110001001110011011.
Note that a paperfolding sequence is 2-automatc if and only if the sequence of unfolding instructions is eventually periodic [7, Theorem 6.5.4].
Allouche [2], and later, Baake [13] proved that no paperfolding sequence contains a palindrome of length >13. In fact, even more is true as shown below.
Proposition 57.
No paperfolding sequence contains a reflected factor of length >13.
Proof.
It suffices to show that no paperfolding sequence contains a reflected factor of length 14. For if this holds, but there is a longer reflected factor x, we could write =xyz where =|y|14. Then =xRzRyR, so y would be a reflected factor of length 14, a contradiction.
Now, by a known result on the appearance function of paperfolding sequences [90, Theorem 12.2.1], we know that every length-14 factor of a paperfolding sequence pf appears in a prefix of length 109, which is in turn specified by the first 7 unfolding instructions. We can then simply examine each of the 56 length-14 factors of these 128 (finite) words and verify that no factor is reflected. ∎
A generalized Golay–Shapiro sequence g is defined by taking the running sum, modulo 2, of a paperfolding sequence pf. The famous Golay–Shapiro sequence (also called the Rudin–Shapiro sequence[56, 57, 86, 91] corresponds to the case of unfolding instructions 0(01)ω [4, Definition 6]. Note that a generalized Golay–Shapiro sequence is 2-automatic if and only if its corresponding generalized paperfolding sequence is 2-automatic.
The following analogue of Proposition 57 can be proved like Proposition 57, by appealing to a known result on the recurrence function of generalized Golay–Shapiro sequences [5, Proposition 4.1], giving that every length-15 factor of a paperfolding sequence pf appears in a prefix of length 2408, which is in turn specified by the first 12 unfolding instructions.
Proposition 58.
No generalized Golay–Shapiro sequence contains a reflected factor of length >14.
We can now prove the following result.
Theorem 59.
Let g be a generalized Golay–Shapiro sequence.
  • (a)
    For all n15, we have rg(n)=ρg(n)=-8n8.
  • (b)
    The reflection complexity of every generalized Golay–Shapiro sequence is the same, and takes the values 2, 3, 6, 10, 14, 22, 30, 42, 48, 62, 72, 83, 92, 103 for n[1,14].
Proof.
We prove each item separately.
  • (a)
    For n15, the result follows from combining the results of Allouche and Bousquet-Melou [5] and Proposition 58.
  • (b)
    The result for n15 follows from Item (a). For <n15 the result can be verified by enumeration of all length-2408 prefixes of paperfolding sequences specified by instructions of length 12.
This ends the proof. ∎
Let the Baum–Sweet sequence
b=(b(n))n0=1101100101001001100100000100100101001001
be defined by =b(0)1 and for n1, b(n) is 1 if the base-2 expansion of n contains no block of successive zeros of odd length and 0 otherwise. It is 2-automatic as well. The factor complexity function for b starts with
=(ρb(n))n01,2,4,7,13,17,21,27,33,38,45,52,59,65,70, (10)
and the reflection complexity function for b with
=(rb(n))n01,2,3,5,8,11,13,17,21,25,30,35,40,46,50,56,. (11)
We can again compute a linear representation for (rb(n))n0 using the following Walnut code:
def factoreq_bs "At (t<n) => BS[i+t]=BS[j+t]"::
def factorreveq_bs "At (t<n) => BS[i+t]=BS[(j+n)-(t+1)]"::
def rc_bs n "Aj (j<i) => ((~\$factoreq_bs(i,j,n))
 & (~\$factorreveq_bs(i,j,n)))"::
This gives us a linear representation of rank 90. From this linear representation, a computation proves the following result.
Corollary 60.
Let b be the Baum–Sweet sequence. Then the first difference of the sequence rb(n) is 2-automatic, over the alphabet {1,2,,8}.
9 Further directions
We conclude the paper by considering some further research directions to pursue in relation to reflection complexities of sequences and by raising some open problems.
We encourage further explorations of the evaluation of rx for sequences x for which properties of Palx and/or ρx are known, especially if Walnut cannot be used directly in the investigation of rx. For example, by letting the Chacon sequence c be the fixed point of the morphism 00010 and 11, it is known that =Palc(n)0 for all n13. Also, its factor complexity satisfies =ρc(n)-2n1 for n2 [49]. We have
=(rc(n))n01,2,2,4,4,6,7,10,11,14,16,20,23,25,27,29,31,33,.
The sequence c is not automatic in a given so-called addable numeration system (i.e., where there is an adder). Therefore, we cannot use Walnut, in this case. However, an inductive argument can be applied to prove that =rc(n)ρc(n) for all n13.
We propose the following natural questions.
Question 61.
To what extent can the reflection complexity be used to discriminate between different families of sequences, by analogy with our characterizations of Sturmian and eventually periodic sequences?
The complexity function Unrx defined above may be of interest in its own right, as is the case with the “reflection-free” complexity function enumerating factors such that the reversal of every sufficiently large factor is not a factor.
Question 62.
How can Theorem 56 be generalized with the use of standard generalizations of the Thue–Morse sequence?
For example, if we let
t3=(t3(n))n0=011212201220200112202001200
denote the generalized Thue–Morse sequence for which the nth term t3(n) is equal to the number of 1’s, modulo 3, in the base-2 expansion of n, it can be shown that =rt3(n)ρt3(n) for all n3, and it appears that a similar property holds for the cases given by taking the number of 1’s modulo >4.
Question 63.
What is the reflection complexity of the Thue–Morse sequence over polynomial extractions, with regard to the work of Moshe [73]?
Question 64.
How can the upper bound in Theorem 21 be improved? If rx(n) is of the form Ω(n), then how can this be improved?
Question 65.
How does the reflection complexity compare with other complexity functions, e.g., the complexity functions listed in Section 1?
This leads us to ask about the respective growths of the complexity functions listed in Section 1, in particular for morphic sequences. In this direction, recall that the factor complexity of a pure morphic sequence is either Θ(1), Θ(n), Θ(nloglogn), Θ(nlogn) or Θ(n2), see [76] (more details can be found, e.g., in [38]). The factor complexity of a morphic sequence is either Θ(n+1/1k) for some positive integer k, or else it is O(nlogn), see [45, 44]). As an illustration with a result that has not been already cited above, a comparison between growths for the factor complexity and the Lempel-Ziv complexity can be found in [39]. We end with an easy result for the growth of the reflection complexity in the case of pure morphic and morphic sequences.
Proposition 66.
  • The reflection complexity of a pure morphic sequence is either Θ(1), Θ(n), Θ(nloglogn), Θ(nlogn) or Θ(n2).
  • The reflection complexity of a morphic sequence is either Θ(n+1/1k) for some positive integer k, or else it is O(nlogn).
Proof.
Use the inequalities in Theorem 9: for a sequence x and for all n0, we have 12ρx(n)rx(n)ρx(n). ∎
Acknowledgments
We thank Boris Adamczewski for interesting discussions. John Campbell is grateful to acknowledge support from a Killam Postdoctoral Fellowship from the Killam Trusts, and thanks Karl Dilcher for a useful discussion. The research of Jeffrey Shallit is supported by NSERC grant 2024-03725. Manon Stipulanti is an FNRS Research Associate supported by the Research grant 1.C.104.24F. The authors are thankful for a number of detailed reviewer reports that have substantially improved this paper.